OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). x - 2y = 0 4x - 3y = 15

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

you do it the same way as the other: get one of the equations in form like x=... or y=... in this case x=2y would be the easiest. then put this new equation into the second one: 4*(2y)-3y=15 8y-3y=15 5y=15 y=3 now you go back to your first equation: x=2y since we now know that y=3 x=2*3 x=6

OpenStudy (anonymous):

in the first one move -2y to the other side making x = 2y substitute 2y in for x 8y - 3y = 15 5y=15 y=3 substitute y into the first or second x-6=0 x=6

OpenStudy (anonymous):

from equation 1, x=2y ...(1) from equation 2, y=(4x-15)/3 ...(2) (1) to (2) by substitution we get y=3 AND X=6